# Fedor V. Fomin

### v2

@@ 3.3 Proof of Theorem 1
Proof of Claim 7. First note that at least $m-4k$ complete connected components are guaranteed in $G-S$ by 6 ($S$ separates $m-2k$ special paths) and inapplicability of 2 (at most $2k$ vertex sets of special paths are not cliques).
@@
Set $\mathcal{S}$ is obtained from a path cover $\mathcal{S^{\prime}}$ of $G^{\prime}$ via adding paths covering $V(G)\setminus V(G^{\prime})$. Hence it is also a path cover. On the other hand, $I$ is (if obtained) an independent set in $G$ because $I\subseteq V(G-S)$ and $I$ is independent in $G-S$.
@@
$|I|-|\mathcal{S}|=2k+(|I|-|I^{\prime}|)-(|\mathcal{S}|-|\mathcal{S}^{\prime}|)>2k+(-k+h-|X|)+(h-|X|)=k.$
@@ 4 Proof of Theorem 2: List TM-embedding parameterized by independence number
In this section, we prove Theorem 2. The proof consists of three main steps, deferred to their respective subsections. The first step is the proof of the stronger version of the theorem for highly connected graphs which we call Spanning Lemma. It shows that for any fixed injection $f:V(H)\to V(G)$ the desired TM-embedding respecting $f$ or an independent set can be found; moreover, the found TM-embedding spans $V(G)$. The second step is the so-called Merging Lemma, that combines the solutions from several highly-connected components, provided that the size of the separator between them is small. Finally, to obtain the proof of Theorem 2 itself, we show an iterative procedure that either finds the desired separator, or an independent set.
@@ 4.1 Spanning Lemma
We first deal with highly-connected graphs. More precisely, we demonstrate that, given an injective mapping $f:V(H)\to V(G)$, we can either compute a TM-embedding of $H$ in $G$ respecting $f$ and spanning $V(G)$, or find an independent set of size $k$ in $G$; this holds if the connectivity of $G$ is sufficiently high. Recall that a graph $G$ is $\ell$-linked if it has at least $2\ell$ vertices and for every set of $\ell$ disjoint pairs $(s_{j},t_{j})$ of distinct vertices, there exist $\ell$ vertex-disjoint paths in $G$ such that the endpoints of the $j$-path are $s_{j}$ and $t_{j}$, for each $j\in[\ell]$. A graph is called $-\ell$-linked if, in addition, the set of $\ell$ paths covers all vertices of $G$.
We need a procedure constructing the linkage for computing a TM-embedding of $H$. For that we need a constructive version of the following theorem of Thomas and Wollan [TW05], which proof is non-constructive.
Proposition 2 (Corollary 1.3, [TW05]).
@@
Let $T=\bigcup_{i=1}^{\ell}\{s_{i},t_{i}\}$ be the set of vertices in the input pairs. We call the vertices in $T$ terminals. For simplicity, our algorithm first ensures that $s_{1},t_{1},s_{2},t_{2},\ldots,s_{\ell},t_{\ell}$ are $2\ell$ distinct vertices, i.e., $|T|=2\ell$. For each vertex $v\in T$ that appears in $p>1$ input pairs, the algorithm replaces $v$ by $p$ true twins, i.e., vertices that contain the neighborhood of $v$ and are also adjacent to each other. Then, each repeating occurrence of $v$ in the input pairs is replaced by a distinct copy. This transformation results in $2\ell$ distinct terminal vertices. The requirement on $10\ell$-connectivity of the graph still holds, and the modification does not increase the maximum size of an independent set. Furthermore, given an independent set in the transformed graph, replacing each copy of a terminal vertex with the corresponding original vertex of $G$ gives an independent set in the original graph. The same is true for paths avoiding terminal vertices: a path between copies is equivalent to the path between original vertices. Therefore, the transformed input is equivalent to the original input, and the respective solutions can be transformed in polynomial time. Henceforth, the algorithm works with the transformed input, meaning that all vertices in the input pairs are distinct.
@@
Let $C$ be a given clique of size $2\ell$ in $G$. Then a family of internally vertex-disjoint paths $P_{1},P_{2},\ldots,P_{\ell}$, where for each $i\in[\ell]$, $P_{i}$ connects $s_{i}$ with $t_{i}$, can be constructed in polynomial time.
Proof of Claim 8. Since $G$ is $2\ell$-connected, by Menger’s theorem [Men27, Die17], there are $2\ell$ internally vertex-disjoint paths connecting vertices of $T=\bigcup_{i=1}^{\ell}\{s_{i},t_{i}\}$ with distinct vertices of $C$. These paths can be constructed in the standard fashion via a network flow algorithm; in particular, using the Ford-Fulkerson algorithm [FF56] to find $\ell$-flow in $\mathcal{O}(\ell\cdot|G|)$ running time. Then for each $i\in[\ell]$, we have two paths, one connecting $s_{i}$ with $c_{2i-1}$ and the other connecting $t_{i}$ with $c_{2i}$ for some $c_{2i-1},c_{2i}\in C$. Note that it is possible that one or two of these paths are trivial, i.e., $s_{i}=c_{2i-1}$ or $t_{i}=c_{2i}$ can be true.
@@
It is left to explain how the algorithm constructs the clique $C$, given that $G$ has at least $(k+2\ell)^{k}$ vertices. The essential tool here is the classical Ramsey’s theorem [Ram30], which guarantees a large clique or independent set in a large enough graph. For completeness, we present the corresponding polynomial-time subroutine based on the Ramsey number bound given by Erdős and Szekeres in 1935 [ES09].
@@
Proof of Claim 10. Note that $G$ necessary has either an independent set of size $r$ or a clique of size $s$ because the corresponding Ramsey number satisfies $R(r,s)\leq\binom{r+s-2}{r-1}$ [ES09]. The algorithm is recursive. Since $r,s>0$, $|V(G)|>1$ and $G$ has at least one vertex, denote it by $v$. If $r=1$ or $s=1$, then algorithm outputs $\{v\}$ and stops. For $r,s\geq 2$, consider vertex sets $A=N_{G}(v)$ and $B=V(G)\setminus\{v\}\setminus A$. Since $|V(G)|=|A|+|B|+1$ and $V(G)\geq\binom{r+s-2}{r-1}$ we have that either $|A|\geq\binom{r+s-3}{r-1}$ or $|B|\geq\binom{r+s-3}{r-2}$. If $|A|\geq\binom{r+s-3}{r-1}$ then $G[A]$ has either an independent set of size $r$ or a clique of size $s-1$. In this case, the algorithm makes a recursive call to itself for the input $G[A]$, $r>0$ and $s-1>0$. If its output is a clique of size $s-1$, the algorithm additionally appends $v$ to it, resulting in a clique of size $s$. Thus, the algorithm obtains either an independent set of size $r$, or a clique of size $s$ as required. The case of $|B|\geq\binom{r+s-3}{r-2}$ is symmetric, as $v$ can be appended to any independent set in $G[B]$.
@@
In order to proceed, it is convenient to reformulate the problem in terms of disjoint paths: Given $\ell$ vertex pairs $(s_{1},t_{1}),(s_{2},t_{2}),\ldots,(s_{\ell},t_{\ell})$ of distinct vertices in $G$, find a family of internally-disjoint paths $P_{1},P_{2},\ldots,P_{\ell}$ such that $P_{i}$ connects $s_{i}$ with $t_{i}$ for each $i\in[\ell]$ and each vertex of $G$ belongs to at least one path, and the union of the paths spans $V(G)$.
@@
We describe the enlargement process, akin to the method outlined in [JK23], with a key distinction. Unlike the scenario in [JK23], we lack assurance that $G$ satisfies $\alpha(G)<k$, making the procedure potentially prone to failure. Therefore, we ensure that (a) the enlargement is completed within polynomial time and (b) if the enlargement fails, an independent set of size $k$ is identified in $G$.
The procedure works as follows: If $P_{1},P_{2},\ldots,P_{\ell}$ span all vertices of $G$, the algorithm achieves the desired outcome, reports the solution, and halts. Otherwise, let $S=\bigcup_{i=1}^{\ell}V(P_{i})$ denote the span of the path sequence, with the assumption that $S\neq V(G)$. The algorithm selects any $x\in V(G)\setminus S$. By the $(k+2)\ell$-connectivity of $G$, Menger’s theorem guarantees the existence of $\min\{(k+2)\ell,|S|\}$ paths. Each path goes from $x$ to $S$, and no two paths share any common vertex (including endpoints in $S$) but $x$. One can find such paths by utilizing the Ford-Fulkerson algorithm [FF56] to find a flow of size at most $(k+2)\ell$ in $\mathcal{O}(k\ell\cdot|G|)$ running time.
@@
It is left for the algorithm to process the case $(k+2)\ell<|S|$. There are precisely $(k+2)\ell$ paths that were found by the Ford-Fulkerson algorithm, so at least $k\ell$ paths end in internal vertices of $P_{1},P_{2},\ldots,P_{\ell}$. Hence, there is $i\in[\ell]$ such that at least $k$ paths end in internal vertices of $P_{i}$. The algorithm takes arbitrary $k$ of them. Denote the $k$ selected paths by $Q_{1},Q_{2},\ldots,Q_{k}$ and their respective endpoints in $P_{i}$ by $v_{1},v_{2},\ldots,v_{k}$. For each $a\in[k]$, let $u_{a}$ be the successor of $v_{a}$ in $P_{i}$ if one follows vertices on $P_{i}$ from $s_{i}$ to $t_{i}$.
@@ 4.2 Merging Lemma
Let $G$ be a graph, let $S\subseteq V(G)$ be a vertex subset in $G$, and let $C_{1}$, …, $C_{t}$ be the connected components of $G-S$. Let $H$ be a graph, let $k$ be a parameter, and $L:V(H)\to 2^{V(G)}$ be a list assignment. Assume that each component $C_{i}$ is at least $\max\{k+2,10\}\cdot(3h+3s)$-connected, where $h=|V(H)|+|E(H)|$, $s=|S|$. There is an algorithm that either returns a maximum-size list TM-embedding of $H$ in $G$ respecting the list assignment $L$, or correctly reports that none exists, or returns an independent set in $G$ of size $k$. The running time of the algorithm is $2^{(s+h+k)^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}$.
@@
Formally, a cut descriptor is the tuple $(W,M_{W},\xi,f_{W})$, where $W$ is a set of vertices (indexed implicitly from $1$ to $|W|$), $M_{W}$ is a graph with the vertex set $S^{\prime}\cup W$ with $S^{\prime}\subseteq S$, $\xi:W\to[t]$ is a mapping from $W$ to indices of the components $C_{1}$, …, $C_{t}$, and $f_{W}:V(H)\to S^{\prime}\cup W$ is an injective mapping from $V(H)$ to vertices of $M_{W}$. Denote $T_{W}=\operatorname{Im}f_{W}$, and for each $i\in[t]$, denote $T_{i}=\{w\in W:\xi(w)\in C_{i}\}$. We require the following properties for $(W,M_{W},\xi,f_{W})$ to be a cut descriptor:
@@ 5. CD5
Proof of Claim 11. We call a vertex $w\in W$ a terminal if $w\in T_{W}$, and a connector otherwise. From CD2, every connector in $w$ is adjacent to a vertex of $S$ in $M_{W}$. Since $f_{W}$ is injective, number of terminals in $W$ is at most $|V(H)|$, therefore it remains to bound the number of connectors in $W$ by $|E(H)|+2s$. Let $w\in W$ be a connector that is adjacent to a terminal in $S$, the number of such connectors is at most $\sum_{s\in S\cap T_{W}}|N_{M_{W}}(s)\cap W|\leq|E(H)|$, since by CD1 every $s\in S\cap T_{W}$ has degree $\deg_{H}(h)$, where $h$ is the unique pre-image of $s$ with respect to $f_{W}$, and no edge of $H$ is counted twice. Finally, the number of connectors in $W$ adjacent only to non-terminals of $S$ is at most $2s$, since every non-terminal has degree two in $M_{W}$ by CD1. ∎
@@
Proof of Claim 12. We construct the desired cut descriptor as follows. First, let $W$ be the subset of vertices in $G$ that are either terminals in $V(G)\setminus S$ or neighbors of $S$ in the model $M$, i.e., $W=(\operatorname{Im}f\setminus S)\cup N_{M}(S)$. We then dissolve all vertices in $V(M)\setminus(S\cup W)$, which are all non-terminals by definition of $W$, to obtain $M_{W}$ from $M$. We set $f_{W}$ as the respective mapping of the terminals obtained from $f$. Clearly, CD1 is fulfilled for $(M_{W},T_{W})$, and CD2 also holds immediately by construction. We set $\xi$ so that $T_{i}=W\cap C_{i}$ for each $i\in[t]$. We now proceed to verify the remaining properties of a cut descriptor.
@@
Finally, for CD5, the assigment $\tau$ is given directly by construction as an identity assignment on $W\subset V(G)$. For each $w\in W$, $\tau(w)\in C_{\xi(w)}$ by the construction of $\xi$. If $f_{W}(h)=w$, then $w=\tau(w)\in L(h)$ since $f(h)=w$ and $(M,f)$ is a list TM-embedding respecting $L$. By construction of $M_{W}$, the edges between $W$ and $S$ are never changed between $M$ and $M_{W}$, therefore $N_{M_{W}}(w)\cap S=N_{M}(w)\cap S$, and the latter is a subset of $N_{G}(w)$ since $M$ is a subgraph of $G$.

### v3

@@ 3.3 Proof of Theorem 1
Proof of Claim 7. First note that at least m4km-4k complete connected components are guaranteed in GSG-S by 6 ($S$ separates m2km-2k special paths) and inapplicability of 2 (at most $2k$ vertex sets of special paths are not cliques).
@@
Set $\mathcal{S}$ is obtained from a path cover $\mathcal{S^{\prime}}$ of $G^{\prime}$ via adding paths covering $V(G)\setminus V(G^{\prime})$. Hence it is also a path cover. On the other hand, $I$ is (if obtained) an independent set in $G$ because $I\subseteq V(G-S)$ and $I$ is independent in GSG-S.
@@
$|I|-|\mathcal{S}|=2k+(|I|-|I^{\prime}|)-(|\mathcal{S}|-|\mathcal{S}^{\prime}|)>2k+(-k+h-|X|)-(h-|X|)=k.$
@@ 4 Proof of Theorem 2: List TM-embedding parameterized by independence number
In this section, we prove Theorem 2. The proof consists of three main steps, deferred to their respective subsections. The first step is the proof of the stronger version of the theorem for highly connected graphs, which we call Spanning Lemma. It shows that for any fixed injection $f:V(H)\to V(G)$, the desired TM-embedding respecting $f$ or an independent set can be found; moreover, the found TM-embedding spans $V(G)$. The second step is the so-called Merging Lemma, which combines the solutions from several highly-connected components, provided that the size of the separator between them is small. Finally, to obtain the proof of Theorem 2 itself, we show an iterative procedure that either finds the desired separator or an independent set.
@@ 4.1 Spanning Lemma
We first deal with highly-connected graphs. More precisely, we demonstrate that, given an injective mapping $f:V(H)\to V(G)$, we can either compute a TM-embedding of $H$ in $G$ respecting $f$ and spanning $V(G)$, or find an independent set of size $k$ in $G$; this holds if the connectivity of $G$ is sufficiently high. Recall that a graph $G$ is $\ell$-linked if it has at least $2\ell$ vertices and for every set of $\ell$ disjoint pairs $(s_{j},t_{j})$ of distinct vertices, there exist $\ell$ vertex-disjoint paths in $G$ such that the endpoints of the $j$-path are $s_{j}$ and $t_{j}$, for each $j\in[\ell]$. A graph is called Hamiltonian-$\ell$-linked if, in addition, the set of $\ell$ paths covers all vertices of $G$.
We need a procedure constructing the linkage for computing a TM-embedding of $H$. For that we need a constructive version of the following theorem of Thomas and Wollan [87] whose original proof is non-constructive.
Proposition 2 (Corollary 1.3, [87]).
@@
Let $T=\bigcup_{i=1}^{\ell}\{s_{i},t_{i}\}$ be the set of vertices in the input pairs. We call the vertices in $T$ terminals. For simplicity, our algorithm first ensures that $s_{1},t_{1},s_{2},t_{2},\ldots,s_{\ell},t_{\ell}$ are $2\ell$ distinct vertices, i.e., $|T|=2\ell$. For each vertex $v\in T$ that appears in $p>1$ input pairs, the algorithm replaces $v$ by $p$ true twins, i.e., by vertices with the same neighborhood as $v$ that are also adjacent to each other. Then, each repeating occurrence of $v$ in the input pairs is replaced by a distinct copy. This transformation results in $2\ell$ distinct terminal vertices. The requirement on $10\ell$-connectivity of the graph still holds, and the modification does not increase the maximum size of an independent set. Furthermore, given an independent set in the transformed graph, replacing each copy of a terminal vertex with the corresponding original vertex of $G$ gives an independent set in the original graph. The same is true for paths avoiding terminal vertices: a path between copies is equivalent to the path between original vertices. Therefore, the transformed input is equivalent to the original input, and the respective solutions can be transformed in polynomial time. Henceforth, the algorithm works with the transformed input, meaning that all vertices in the input pairs are distinct.
@@
Let $C$ be a given clique of size $2\ell$ in $G$. Then a family of vertex-disjoint paths $P_{1},P_{2},\ldots,P_{\ell}$, where for each $i\in[\ell]$, $P_{i}$ connects $s_{i}$ with $t_{i}$, can be constructed in polynomial time.
Proof of Claim 8. Since $G$ is $2\ell$-connected, by Menger’s theorem [78, 28], there are $2\ell$ internally vertex-disjoint paths connecting vertices of $T=\bigcup_{i=1}^{\ell}\{s_{i},t_{i}\}$ with distinct vertices of $C$. These paths can be constructed in the standard fashion via a network flow algorithm; in particular, using the Ford-Fulkerson algorithm [42] to find $\ell$-flow in $\mathcal{O}(\ell\cdot|G|)$ running time. Then for each $i\in[\ell]$, we have two paths, one connecting $s_{i}$ with $c_{2i-1}$ and the other connecting $t_{i}$ with $c_{2i}$ for some $c_{2i-1},c_{2i}\in C$. Note that it is possible that one or two of these paths are trivial, i.e., $s_{i}=c_{2i-1}$ or $t_{i}=c_{2i}$ can be true.
@@
It is left to explain how the algorithm constructs the clique $C$, given that $G$ has at least $(k+2\ell)^{k}$ vertices. The essential tool here is the classical Ramsey’s theorem [84], which guarantees a large clique or independent set in a large enough graph. For completeness, we present the corresponding folklore polynomial-time subroutine based on the Ramsey number bound given by Erdős and Szekeres in 1935 [32].
@@
Proof of Claim 10. Note that $G$ necessarily has either an independent set of size $r$ or a clique of size $s$ because the corresponding Ramsey number satisfies $R(r,s)\leq\binom{r+s-2}{r-1}$ [32]. The algorithm is recursive. Since $r,s>0$, $|V(G)|\geq 1$ and $G$ has at least one vertex, denote it by $v$. If $r=1$ or $s=1$, then algorithm outputs $\{v\}$ and stops. For $r,s\geq 2$, consider vertex sets $A=N_{G}(v)$ and $B=V(G)\setminus\{v\}\setminus A$. Since $|V(G)|=|A|+|B|+1$ and $V(G)\geq\binom{r+s-2}{r-1}$ we have that either $|A|\geq\binom{r+s-3}{r-1}$ or $|B|\geq\binom{r+s-3}{r-2}$. If $|A|\geq\binom{r+s-3}{r-1}$ then $G[A]$ has either an independent set of size $r$ or a clique of size s1s-1. In this case, the algorithm makes a recursive call to itself for the input $G[A]$, $r>0$ and s1>0s-1>0. If its output is a clique of size s1s-1, the algorithm additionally appends $v$ to it, resulting in a clique of size $s$. Thus, the algorithm obtains either an independent set of size $r$, or a clique of size $s$ as required. The case of $|B|\geq\binom{r+s-3}{r-2}$ is symmetric, as $v$ can be appended to any independent set in $G[B]$.
@@
In order to proceed, it is convenient to reformulate the problem in terms of disjoint paths: Given $\ell$ pairs $(s_{1},t_{1}),(s_{2},t_{2}),\ldots,(s_{\ell},t_{\ell})$ of distinct vertices in $G$, find a family of internally-disjoint paths $P_{1},P_{2},\ldots,P_{\ell}$ such that $P_{i}$ connects $s_{i}$ with $t_{i}$ for each $i\in[\ell]$ and each vertex of $G$ belongs to at least one path, and the union of the paths spans $V(G)$.
@@
We describe the enlargement process, akin to the method outlined in [61], with a key distinction. Unlike the scenario in [61], we lack assurance that $G$ satisfies $\alpha(G)<k$, making the procedure potentially prone to failure. Therefore, we ensure that (a) the enlargement is completed within polynomial time and (b) if the enlargement fails, an independent set of size $k$ is identified in $G$.
The procedure works as follows: If $P_{1},P_{2},\ldots,P_{\ell}$ span all vertices of $G$, the algorithm achieves the desired outcome, reports the solution, and halts. Otherwise, let $S=\bigcup_{i=1}^{\ell}V(P_{i})$ denote the span of the path sequence, with the assumption that $S\neq V(G)$. The algorithm selects any $x\in V(G)\setminus S$. By the $(k+2)\ell$-connectivity of $G$, Menger’s theorem guarantees the existence of $\min\{(k+2)\ell,|S|\}$ paths. Each path goes from $x$ to $S$, and no two paths share any common vertex (including endpoints in $S$) but $x$. One can find such paths by utilizing the Ford-Fulkerson algorithm [42] to find a flow of size at most $(k+2)\ell$ in $\mathcal{O}(k\ell\cdot|G|)$ running time.
@@
It is left for the algorithm to process the case $(k+2)\ell<|S|$. There are precisely $(k+2)\ell$ paths that were found by the Ford-Fulkerson algorithm, so at least $k\ell$ paths end in internal vertices of $P_{1},P_{2},\ldots,P_{\ell}$. Hence, there is $i\in[\ell]$ such that at least $k$ paths end in internal vertices of $P_{i}$. The algorithm arbitrary takes $k$ such paths. Denote the $k$ selected paths by $Q_{1},Q_{2},\ldots,Q_{k}$ and their respective endpoints in $P_{i}$ by $v_{1},v_{2},\ldots,v_{k}$. For each $a\in[k]$, let $u_{a}$ be the successor of $v_{a}$ in $P_{i}$ if one follows vertices on $P_{i}$ from $s_{i}$ to $t_{i}$.
@@
Finally, we note that Lemma 4 admits an existential variant as a direct corollary: If $G$ is sufficiently well-connected compared to its independence number $\alpha(G)$ and the size of $H$, then a spanning TM-embedding of $H$ exists in $G$ for any choice of $f$. As this may be of independent interest, we state formally the following.
Proposition 3.
Let $H$ and $G$ be non-empty graphs, and let $f\colon V(H)\rightarrow V(G)$ be an injective mapping. Let $G$ additionally be $(\max\{\alpha(G)+3,10\}\cdot h)$-connected, where $h=|V(H)|+|E(H)|$. There exists a subgraph $M$ of $G$ such that $(M,f)$ is a TM-embedding of $H$ in $G$ spanning $V(G)$.
@@ 4.2 Merging Lemma
Let $G$ be a graph, let $S\subseteq V(G)$ be a vertex subset in $G$, and let $C_{1}$, …, $C_{t}$ be the connected components of GSG-S. Let $H$ be a graph, let $k$ be a parameter, and $L:V(H)\to 2^{V(G)}$ be a list assignment. Assume that each component $C_{i}$ is at least $\max\{k+2,10\}\cdot(3h+3s)$-connected, where $h=|V(H)|+|E(H)|$, $s=|S|$. There is an algorithm that either returns a maximum-size list TM-embedding of $H$ in $G$ respecting the list assignment $L$, or correctly reports that none exists, or returns an independent set in $G$ of size $k$. The running time of the algorithm is $2^{(s+h+k)^{\mathcal{O}(k)}}\cdot|G|^{\mathcal{O}(1)}$.
@@
Formally, a cut descriptor is the tuple $(W,M_{W},\xi,f_{W})$, where $W$ is a set of vertices (indexed implicitly from $1$ to $|W|$), $M_{W}$ is a graph with the vertex set $S^{\prime}\cup W$ with $S^{\prime}\subseteq S$, $\xi:W\to[t]$ is a mapping from $W$ to indices of the components $C_{1}$, …, $C_{t}$, and $f_{W}:V(H)\to S^{\prime}\cup W$ is an injective mapping from $V(H)$ to vertices of $M_{W}$. Denote $T_{W}=\operatorname{Im}f_{W}$, and for each $i\in[t]$, denote $T_{i}=\{w\in W:\xi(w)=i\}$. We require the following properties for $(W,M_{W},\xi,f_{W})$ to be a cut descriptor:
@@ 5. CD5
Proof of Claim 11. We call a vertex $w\in W$ a terminal if $w\in T_{W}$, and a connector otherwise. From CD2, every connector in $w$ is adjacent to a vertex of $S$ in $M_{W}$. Since $f_{W}$ is injective, the number of terminals in $W$ is at most $|V(H)|$, therefore it remains to bound the number of connectors in $W$ by $|E(H)|+2s$. Let $w\in W$ be a connector that is adjacent to a terminal in $S$, the number of such connectors is at most $\sum_{s\in S\cap T_{W}}|N_{M_{W}}(s)\cap W|\leq|E(H)|$, since by CD1 every $s\in S\cap T_{W}$ has degree $\deg_{H}(h)$, where $h$ is the unique pre-image of $s$ with respect to $f_{W}$, and no edge of $H$ is counted twice. Finally, the number of connectors in $W$ adjacent only to non-terminals of $S$ is at most $2s$, since every non-terminal has degree two in $M_{W}$ by CD1. ∎
@@
Proof of Claim 12. We construct the desired cut descriptor as follows. First, let $W$ be the subset of vertices in $G$ that are either terminals in $V(G)\setminus S$ or neighbors of $S$ in the model $M$, i.e., $W=(\operatorname{Im}f\setminus S)\cup N_{M}(S)$. We then dissolve all vertices in $V(M)\setminus(S\cup W)$, which are all non-terminals by definition of $W$, to obtain $M_{W}$ from $M$; here the dissolving a vertex $x$ of degree two with the neighbors $y$ and $z$ deletes $x$ and adds the edge $yz$. We set $f_{W}$ as the respective mapping of the terminals obtained from $f$. Clearly, CD1 is fulfilled for $(M_{W},T_{W})$, and CD2 also holds immediately by construction. We set $\xi$ so that $T_{i}=W\cap C_{i}$ for each $i\in[t]$. We now proceed to verify the remaining properties of a cut descriptor.
@@
Finally, for CD5, the assignment $\tau$ is given directly by construction as an identity assignment on $W\subset V(G)$. For each $w\in W$, $\tau(w)\in C_{\xi(w)}$ by the construction of $\xi$. If $f_{W}(h)=w$, then $w=\tau(w)\in L(h)$ since $f(h)=w$ and $(M,f)$ is a list TM-embedding respecting $L$. By construction of $M_{W}$, the edges between $W$ and $S$ are never changed between $M$ and $M_{W}$, therefore $N_{M_{W}}(w)\cap S=N_{M}(w)\cap S$, and the latter is a subset of $N_{G}(w)$ since $M$ is a subgraph of $G$.